OpenStudy (rebeccapink6398):

@31356 :D . In a fruit basket containing apples and oranges, 1/6 of the fruit are oranges. Which could be the numbers of oranges and apples in the basket

OpenStudy (texaschic101):

hello texas girls

OpenStudy (rebeccapink6398):

HELLO! :D :D :D

OpenStudy (31356):

Hey, lol :D

OpenStudy (texaschic101):

ok...your problem... 1/6 of the fruit are oranges....that means 5/6 of the fruit are apples. Is this a multiple choice question ?

OpenStudy (31356):

Yea, that was what I was going to type too, lol :D

OpenStudy (rebeccapink6398):

no just one.. A. 3 oranges and 15 apples B. 1 orange and 6 apples C. 15 oranges and 3 apples D. 6 oranges and 1 apple

OpenStudy (texaschic101):

3 oranges and 15 apples (answer A) there are a total of 18 pieces of fruit... 1/6 are oranges... 1/6 (18) = 18/6 = 3 oranges 5/6 are apples 5/6(18) = 90/6 = 15 apples

OpenStudy (31356):

Oh sorry, I wasn't here for a while..

OpenStudy (31356):

Yea, I was going to type that too, lol :D @texaschic101 is correct! :D

OpenStudy (texaschic101):

:) confirmation is always good

OpenStudy (31356):

Triple smileys :D:D:D

OpenStudy (texaschic101):

I am doing the happy dance right now :)

OpenStudy (31356):

Lol 10^10000 times :D

OpenStudy (texaschic101):

time to find more questions....c yall later....thats texas talk..lol

OpenStudy (31356):

Yea, me too :D

OpenStudy (31356):

Thanks for the medal! :D

OpenStudy (rebeccapink6398):

:D...awe :( I got 60% D;

OpenStudy (31356):

How come?? :(

OpenStudy (rebeccapink6398):

not sure P( if gonna retake it

OpenStudy (rebeccapink6398):

:(*